cadedog55:

The Frostburg-Truth bus travels on a straight road from Frostburg Mall to Sojourner Truth Park. The mall is 3 miles west and 4 miles south of the City Center. The park is 3 miles east and 5 miles north of the Center. How far is it from the mall to the park to the nearest tenth of a mile? 1.0 mile 5.0 miles 5.8 miles 10.8 miles

ZoeyBiocth16:

the best answer to this is We travel from mall to park in a straight line. The east-west distance is 3-(-3) miles, or 6 miles. The north-south distance is 5+4, or 9 miles. Use the Pyth. Thm. to find the distance between mall and park (6 miles)^2 + (9 miles)^2 = distance^2 distance = sqrt (36 + 81) = sqrt(117 miles^2) = 10.8 miles.

ZoeyBiocth16:

@cadedog55 do you understand it now